OpenStudy (anonymous):

Which of the following is a solution of x^2+5x=-2? -2+2i√3 2+2i√3 1-i√3 -1+i√3

Nnesha (nnesha):

quadratic equation is Ax^2 + Bx+C=0 so first move -2 to the left side and then use quadratic formula

OpenStudy (anonymous):

use the quadratic formula, do you know that?

OpenStudy (anonymous):

I don't remember the formula

OpenStudy (anonymous):

x = -b +/- sqrt (a^2 - 4ac) all over 2a

OpenStudy (anonymous):

I put the wrong question to those answers. The question is x^2+2x+4

Nnesha (nnesha):

:)\[\huge\ x = \frac{ -b \pm \sqrt{b^2 -4ac} }{ 2a }\]

Nnesha (nnesha):

okay for that one just find two number if you multiply them you should get 4 and if you add or subtract them you should get 2

OpenStudy (anonymous):

once you plug in - A is the correct answer. I uploaded the pix also...caught I typed a^2 instead of b^2

OpenStudy (anonymous):

no, still have to use quad formula.

Nnesha (nnesha):